George is 20 years younger than Jane. 2 years ago, Jane's age was 2 times George's age. How old is George now?
Answer by ikleyn(53763)

.
George is 20 years younger than Jane. 2 years ago, Jane's age was 2 times George's age.
How old is George now?
~~~~~~~~~~~~~~~~~~~

Let x be the George's age.

Then Jane's age is (x+20) years.


2 years ago

    (x+20)-2 = 2*(x-2).


Simplify and find x

    x + 18 = 2x - 4

    18 + 4 = 2x - x

      22   =    x.


ANSWER.  George is 22 years old.  Jane is 22+20 = 42 years old.

Solved.